OpenStudy (anonymous):

Which set represents a Pythagorean Triangle? A: 1,2,4 B: 2,4,6 C: 3,4,10 D: 5,12,13

OpenStudy (anonymous):

well a pythagorean triangle is a right angle one, and its asquared+bsquared=csquared

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

a pythagorean triangle is always a right triangle (it's angle is 90 degrees) and the thereom is \[a^2+b^2=c^2\] That should help you solve it

OpenStudy (anonymous):

the csquared is the hypotenuse, which is the longest side of a triangle (the slanted)

OpenStudy (anonymous):

i'm still confused i'm sorry :( @BeckieBird

OpenStudy (anonymous):

:/ I think the answer is C but I'm not sure it's been a while since I've done these

OpenStudy (anonymous):

tht's wht i thought but idk so thank ya tho

OpenStudy (anonymous):

think of it like this; the legs or first two numbers, then add them and if you get the third number, or hypotenuse, then you are correct. For example: 59, 41, and 100 is a Pythagorean triple or triangle because 59+41=100. here's a calc: http://www.algebra.com/calculators/geometry/pythagorean.mpl copy and paste.